OpenStudy (anonymous):

a line passes through the points (-10 -4) and (-1 2). what is the slope

OpenStudy (the_fizicx99):

M = Y2-y1/X2-x1 M = 2-(-4)/-1-(-10)

OpenStudy (anonymous):

thank u

OpenStudy (the_fizicx99):

Yw